OpenStudy (anonymous):

A roof rises 2.8 feet vertically and 8.1 feet horizontally. What is the grade of the roof?

myininaya (myininaya):

2.8/8.1=28/81

OpenStudy (anonymous):

what is the percentage?

OpenStudy (anonymous):

2.8/8.1=28/81=34.57 percent